A spring of natural length 10 inches stretches 1.5 inches under a
weight of 8 pounds. Find the work done in stretching the spring.
(a) from its natural length to a length of 14 inches
(b) from a length of 11 inches to a length of 13 inches.
1
Expert's answer
2020-06-09T19:02:28-0400

Convert length to metre and mass into kilogram for simplicity.

natural length, "l_0 = 10in = 0.254m"

stretching, "x = 1.5in = 0.0381m"

weight, "m = 8pounds = 3.632kg"

let spring constant is "k"


then "k = \frac{mg}{x} = \frac{3.632*10}{0.0381} = 953.3 Nm^{-1}"


(i) work done in stretching spring from normal to 14 inches.


"W = \frac{1}{2}k(l_2^{2} - l_1^{2}) = \frac{1}{2}*953.3*((0.3556)^{2} - (0.254)^{2}) = 29.5 J"


(ii) work done in stretching from 11 inches to 13 inches


"W = \frac{1}{2}k(l_2^{2} - l_1^{2}) = \frac{1}{2}*953.3*((0.3302)^{2} - (0.2794)^{2})=14.76J"